### v1

# A Note on an Algebraic Criterion for Compact Kähler Manifolds

A Note on an Algebraic Criterion for Compact Kähler Manifolds
Junyan Cao Université Côte d’Azur, CNRS, LJAD, France junyan.cao@unice.fr
1. Introduction
In a recent paper [Ou25], W. Ou generalized the algebraic integrability criteria of Campana-Păun [CP19], Druel [Dru18] and the uniruledness criterion [BDPP] to the compact Kähler setting. A key ingredient in his proof is an algebraicity criterion [Ou25, Thm 1.3], which extends the algebraicity criteria in [Bos01, BM16]. In this note, we explain a proof of his criterion.
Theorem 1.1.
Let $X$ be a compact Kähler manifold and let $C\subset X$ be a compact submanifold. Let $S_{0}$ be an irreducible locally closed submanifold containing $C_{0}$, where $C_{0}$ is an open set in $C$ with ${\rm codim}_{C}(C\setminus C_{0})\geq 2$. Let $M$ be the Zariski closure of $S_{0}$ in $X$.
If $\dim M>\dim S_{0}$, then the conormal bundle $\mathcal{N}^{*}_{C_{0}/S_{0}}$ is pseudo-effective (”psef” for short), in the sense of Definition 2.4.
The major part of our proof of Theorem 1.1 follows closely the ideas in [Ou25]. The main difference is that [Ou25] requires some algebraic constructions to carefully desingularize the singular locus of $M$ and to control the loss of psef of the conormal bundle in a very efficient way. In our proof, we use an analytic approach by comparing the Lelong numbers of a certain current and its pullback, cf. Proposition 2.2. This makes the proof purely analytic.
Note that [Ou25, Thm 1.3] treats the case when ${\rm codim}_{C}(C\setminus C_{0})=1$ with some additional conditions on $C\setminus C_{0}$. Moreover, as we use the analytic Definition 2.4 for the pseudo-effectiveness of sheaves, Theorem 1.1 is slightly weaker than [Ou25, Thm 1.3]. However, as explained in Remark 2.5, Theorem 1.1 still implies the main geometric applications [Ou25, Thm 1.1, Thm 1.4].
The main idea of the proof of Theorem 1.1 is to as follows. We first assume that $M$ is smooth. By using Demailly’s mass concentration method [Dem93] and Collins-Tosatti’s extension theorem [CT14], we construct an $\omega_{X}$-psh function $\varphi$ on $X$ such that the Lelong number $\nu(\varphi|_{S_{0}},y_{0})$ at a fixed point $y_{0}\in C_{0}$ can be arbitrarily large cf. Proposition 3.1. This is one of the key ideas in [Ou25]. Now thanks to Proposition 3.2, we can show that the $\nu(\varphi|_{S_{0}},C_{0})$ can be arbitrary large. In the general case, when $M$ is not smooth, we need additional arguments. Let $\pi:\widehat{X}\to X$ be the blow up of $C$ in $X$. Let $\widehat{S}_{0}$ be the strict transform of $S_{0}$. In Proposition 3.4, we construct $\omega_{\widehat{X}}$-psh functions $\varphi$ on $\widehat{X}$ and show that the Lelong number of $\varphi|_{\widehat{S}_{0}}$ along $\widehat{S}_{0}\cap\pi^{-1}(C_{0})$ can be arbitrarily large. Once this is established, we complete the proof of the main theorem using standard arguments.
Acknowledgements The author would like to thank Wenhao Ou for sending the preliminary versions of his very nice paper [Ou25], as well as for many discussions related to his paper and this note. He would also like to thank Mihai Păun for reading this note and for his many suggestions on it. He thanks the Institut Universitaire de France and the A.N.R JCJC project Karmapolis (ANR-21-CE40-0010) for providing excellent working condition.
2. Some Preparations
We first recall some basic properties of quasi-plurisubharmonic (quasi-psh) functions and positive currents.
Proposition 2.1.
Let $X$ be a complex manifold, and let $X_{0}\subset X$ be an open set such that ${\rm codim}_{X}(X\setminus X_{0})\geq 2$. Let $Y\subset X$ be a submanifold. Then we have:
  • El Mir extension theorem [Dem, Chapter III, Thm 2.3] Let $T$ be a $d$-closed positive $(1,1)$-current on $X_{0}$. Then $T$ can be extended to a positive current on the entire space $X$.
  • Support theorem [Dem, Chapter III, Thm 2.10] Let $T_{1}$ and $T_{2}$ be two $d$-closed positive $(1,1)$-currents on $X$ such that $T_{1}=T_{2}$ on $X_{0}$. Then $T_{1}=T_{2}$ on $X$.
  • [Hor] Let $0\in\mathbb{C}^{n}$ be the origin. Let $\varphi$ be a psh function in a neighborhood of 00. Then the Lelong number $\nu(\varphi,0)$ can be computed using the following formula:
    (1) $\nu(\varphi,0)=\min\{\nu(\varphi|_{L_{\xi}},0)\mid\xi\in\mathbb{C}^{n}\setminus\{0\}\},$
    where $L_{\xi}$ is the complex line $\mathbb{C}\cdot\xi$. In fact, $\nu(\varphi,0)=\nu(\varphi|_{L_{\xi}},0)$ for almost every $\xi\in\mathbb{C}^{n}\setminus\{0\}$.
    As a consequence, for psh functions with analytic singularities, namely $\varphi=c\log\sum_{i=1}^{m}|f_{i}|^{2}+O(1)$, where $f_{i}$ are holomorphic functions, $\frac{1}{2c}\nu(\varphi,0)$ equal to the minimal vanishing order of $(f_{1},\dots,f_{m})$.
  • Let $\varphi$ be a quasi-psh function on $X$ such that $\varphi$ is not identically $\infty$- ∞ on $Y$. Let $\nu(\varphi,Y)$ be the Lelong number of $\varphi$ along $Y$, defined as the Lelong number of $\varphi$ at a generic point of $Y$. By the semicontinuity of Lelong numbers, we have $\nu(\varphi,Y)=\min_{y\in Y}\nu(\varphi,y)$.
  • Let $\pi:\widehat{X}\to X$ be the blow-up along $Y$. Let $x\in X$. If $\varphi$ is a quasi-psh function on $X$ with analytic singularities, by (1) we have
    $\nu(\varphi,x)=\nu(\pi^{*}\varphi,\pi^{-1}(x)),\quad\text{and}\quad\nu(\varphi,Y)=\nu(\pi^{*}\varphi,\pi^{-1}(Y)).$
  • [Dem12, Prop 5.10] Let $\varphi$ be a psh function on $X$. Let $\mathcal{I}(\varphi)$ be the multiplier ideal sheaf generated locally by holomorphic functions $f$ such that
[…]
    If $\varphi=c\log\sum_{i=1}^{n}|z_{i}|^{2\alpha_{i}}$ for real numbers $\alpha_{i}>0$, then $\mathcal{I}(\varphi)_{0}$ is generated by the monomials $\prod z_{i}^{\beta_{i}}$ such that
    $\sum_{i=1}^{n}\frac{\beta_{i}+1}{\alpha_{i}}>2c.$
We will now discuss the Lelong number of quasi-psh functions after a bimeromorphic transformation.
Lemma 2.2.
Let $X$ be a complex manifold, and let $Y\subset X$ be a submanifold. Let $\pi:\widehat{X}\to X$ be the blow-up of $Y$, and let $E\subset\widehat{X}$ be the exceptional divisor. Let $\widehat{T}\geq 0$ be a positive $(1,1)$-current on $\widehat{X}$ with analytic singularities. Let $\widehat{x}\in E$ and set $x:=\pi(\widehat{x})$. Define $T:=\pi_{*}(\widehat{T})$. Then we have
$C\cdot\nu(T,x)\geq\nu(\widehat{T},\widehat{x})-\nu(\widehat{T},E).$
for some uniform constant $C$ depending only on $X$ and $Y$.
Proof.
Set $T_{1}:=\pi^{*}T$. Then $T_{1}$ is a positive current on $\widehat{X}$ and coincides with $\widehat{T}$ outside $E$. Let $a:=\nu(\widehat{T},E)$ and $b:=\nu(T_{1},E)$. Then $(T_{1}-b[E])-(\widehat{T}-a[E])$ is a current supported on a subvariety of codimension at least $2$. By Proposition 2.1, we have
$T_{1}-b[E]\equiv\widehat{T}-a[E]\qquad\text{ on }\widehat{X}.$
Using [Fav99, Thm 2], we know that $\nu(T_{1},\widehat{y})\leq C\nu(T,y)$ for some constant $C$. Therefore, we obtain
$\nu(\widehat{T},\widehat{x})-a=\nu(\widehat{T}-a[E],\widehat{x})=\nu(T_{1}-b[E],\widehat{x})\leq\nu(T_{1},\widehat{x})\leq C\nu(T,x).$
Lemma 2.3.
Let $X$ be a complex Kähler manifold, and let $V$ be a holomorphic vector bundle of rank $r$ on $X$. Let $\pi:\mathbb{P}(V)\to X$ be the natural projection. Fix a point $x\in X$. Let $\omega$ be a Kähler metric on $\mathbb{P}(V)$.
Let $T\geq 0$ be a $d$-closed positive $(1,1)$-current on $\mathbb{P}(V)$. Consider the $(1,1)$-positive current $T^{\prime}:=\pi_{*}(T\wedge\omega^{r-1})$ on $X$. Then we have
$\nu(T^{\prime},x)\geq(\int_{\pi^{-1}(x)}\omega^{r-1})\cdot\nu(T,\pi^{-1}(x)).$
Proof.
We first assume that $\dim X=1$. Set $a:=\nu(T,\pi^{-1}(x))$. Then $T_{a}:=T-a[\pi^{-1}(x)]$ is a $(1,1)$-positive current. We have
$T^{\prime}=\pi_{*}(T_{a}\wedge\omega^{r-1})+a\left(\int_{\pi^{-1}(x)}\omega^{r-1}\right)[x]\geq a\left(\int_{\pi^{-1}(x)}\omega^{r-1}\right)[x].$
Thus, the lemma is proved.
If $\dim X\geq 1$, we take a generic one-dimensional unit disc $\Delta\subset X$ passing through $x$. By (1), we have $\nu(T^{\prime},x)=\nu(T^{\prime}|_{\Delta},x)$. Thus, the proposition is reduced to the one-dimensional case. ∎
Definition 2.4.
Let $X$ be a compact Kähler manifold, and let $\mathcal{F}$ be a reflexive sheaf on $X$. Let $X_{0}\subset X$ be the locally free locus of $\mathcal{F}$. Consider the projection $\pi:\mathbb{P}(\mathcal{F}|_{X_{0}})\to X_{0}$, and let $Y$ be a desingularization of $\mathbb{P}(\mathcal{F})$ with a Kähler metric $\omega_{Y}$. Then $\mathbb{P}(\mathcal{F}|_{X_{0}})$ can be embedded in $Y$.
We say that $\mathcal{F}|_{X_{0}}$ is psef if, for every $\epsilon>0$, there exists a possibly singular metric $h_{\epsilon}$ such that $i\Theta_{h_{\epsilon}}(\mathcal{O}_{\mathbb{P}(\mathcal{F}|_{X_{0}})}(1))\geq-\epsilon\omega_{Y}$ on $\mathbb{P}(\mathcal{F}|_{X_{0}})$.
Remark 2.5.
Since $\mathbb{P}(\mathcal{F}|_{X_{0}})$ is not compact (its complement may have codimension $1$), it is not clear whether the above definition implies the existence of a metric $h$ such that $i\Theta_{h}(\mathcal{O}_{\mathbb{P}(\mathcal{F}|_{X_{0}})}(1))\geq 0$ on $\mathbb{P}(\mathcal{F}|_{X_{0}})$. Thus, Definition 2.4 is à priori weaker than [Ou25, Definition 4.1] and [HP19].
However, if $\mathcal{F}|_{X_{0}}$ is psef in the sense of Definition 2.4, then by applying the same argument as in [Ou25, Section 4] (replacing $h$ with $h_{\epsilon}$ and letting $\epsilon\to 0$), we can show that $\mu_{\alpha,\max}(\mathcal{F})\geq 0$ for any mobile class $\alpha$. Therefore, Theorem 1.1 still implies the main geometric applications in [Ou25, Thm 1.1, Thm 1.4].
3. Proof of the Main Results
Let $X$ be a compact Kähler manifold, and let $S_{0}$ be a locally closed submanifold of $X$. Fix a point $x\in S_{0}$. When $X$ is projective, fix an ample line bundle $A$ on $X$. Bost [Bos04] proved a characterization of the algebraicity of $S_{0}$ by studying the subspaces $E^{i}_{m}\subset H^{0}(X,mA)$ consisting of sections whose restrictions to $S_{0}$ vanish to order at least $i$ at $x$. In [Ou25], Ou generalized Bost’s nice idea to the setting of compact Kähler manifolds by studying the restricted Lelong number $\nu(\varphi|_{S_{0}},x)$, where $\varphi$ is an $\omega_{X}$-plurisubharmonic function on $X$ for some fixed Kähler metric $\omega_{X}$. For the reader’s convenience, we first recall his elegant proof, which follows precisely the same argument as the first part of [Ou25, Thm 1.2].
Proposition 3.1.
Let $X$ be a compact Kähler manifold, and let $\omega_{X}$ be a Kähler metric on $X$. Let $S_{0}$ be a locally closed submanifold of $X$ such that $\dim S_{0}<\dim X$ and the Zariski closure of $S_{0}$ is $X$. Fix a point $x\in S_{0}$.
Then, for any $\lambda>0$, we can find a $\omega_{X}$-psh function $\varphi$ on $X$ such that $\varphi$ has analytic singularities and satisfies $\nu(\varphi|_{S_{0}},x)\geq\lambda$.
Proof.
Let $U$ be a neighborhood of $x$ in $X$, and let $V$ be a neighborhood of $x$ in $S_{0}$. Since $\dim S_{0}<\dim X$, we can find a local holomorphic coordinate system $(z_{1},\dots,z_{n})$ on $U$ such that $V$ is contained in $\{z_{1}=0\}$.
Let $m\in\mathbb{N}$. We consider the function
$f=\frac{1}{m^{\frac{n-1}{n}}}\log(|z_{1}|^{2}+|z_{2}|^{2m}+\cdots+|z_{n}|^{2m}).$
Since $V$ is contained in $\{z_{1}=0\}$, we have $\nu(f|_{S_{0}},x)=2m^{\frac{1}{n}}$. By construction, the mass $(i\partial\overline{\partial}f)^{n}=A\cdot\{x\}$ for some uniform constant $A$ independent of $m$. By applying Demailly’s mass concentration theorem [Dem93, Cor 6.8], we can find a $2\omega_{X}$-psh function $\psi$ on $X$ such that
$\psi\leq C\cdot f+O(1)$
in a neighborhood of $x$, where $C$ is a uniform constant independent of $m$.
Next, we apply Demailly’s regularization theorem to $\psi$. We can find a $\omega_{X}$-psh function $\varphi$ with analytic singularities, and locally, $\varphi$ is of the form
$\varphi=\frac{1}{k}\log\sum|g_{i}|^{2}+O(1),$
where $\int|g_{i}|^{2}e^{-2k\psi}<+\infty$ for some sufficiently large $k$. Now, the key point is that since $S_{0}$ is dense in $X$ and $\varphi$ is of analytic singularity, $\varphi|_{S_{0}}$ is well defined.
Now we calculate $\nu(\varphi|_{S_{0}},x)$. Since $\psi\leq Cf$, we know that $g_{i}\in\mathcal{I}(kCf)$. By Proposition 2.1, $\mathcal{I}(kCf)$ is generated by monomials of the form $z_{1}^{a_{1}}\cdots z_{n}^{a_{n}}$ such that
$a_{1}+1+\sum_{i=2}^{n}\frac{a_{i}+1}{m}>\frac{kC}{m^{\frac{n-1}{n}}}.$
Thus, for every $g_{i}$, the vanishing order of $(g_{i}|_{z_{1}=0},0)$ is at least
$\sum_{i=2}^{n}a_{i}>m^{\frac{1}{n}}kC-n+1.$
By Proposition 2.1, we obtain
$\nu(\varphi|_{S_{0}},x)\geq m^{\frac{1}{n}}\cdot C-\frac{n-1}{k}.$
The proposition follows by choosing $m$ such that $m^{\frac{1}{n}}\geq\frac{\lambda}{C}$. ∎
Let $C$ be a compact submanifold of $X$ such that there exists a Zariski open subset $C_{0}\subset C$ with ${\rm codim}_{C}(C\setminus C_{0})\geq 2$ and $C_{0}\subset S_{0}$. Let $y_{0}\in C_{0}$. We know that $\nu(\varphi|{S_{0}},y_{0})\geq\nu(\varphi|{S_{0}},C_{0})$. The following proposition provides a control in another direction. The proof partially follows the idea in the second part of [Ou25, Thm 1.2]. The main difference is that, instead of using the algebraic construction in [Ou25, Section 5] to control the singularity of the currents on $\mathbb{P}(\mathcal{N}^{*}_{C_{0}/S_{0}})$ near its boundary, we use Lemma 2.3 to directly obtain a positive current on the compact manifold $C$.
Proposition 3.2.
Let $X$ be a compact Kähler manifold and $C\subset X$ be a compact submanifold. Let $S_{0}$ be a locally closed submanifold containing $C_{0}$, where $C_{0}$ is an open subset of $C$ with ${\rm codim}_{C}(C\setminus C_{0})\geq 2$.
We fix a point $y_{0}\in C_{0}$. Let $\omega_{X}$ be a Kähler metric on $X$. Then there exists some constants $K_{1}$ and $K_{2}$ depending only on $(\omega_{X},y_{0},C_{0},S_{0},X)$ such that, for any $\omega_{X}$-psh function $\varphi$ on $X$, if $\varphi|_{S_{0}}$ is of algebraic singularities, then we have
$\nu(\varphi|_{S_{0}},y_{0})\leq K_{1}\cdot\nu(\varphi|_{S_{0}},C_{0})+K_{2}.$
Proof.
Let $\pi:\widehat{X}\to X$ be the blow-up of $C$ in $X$. Let $\widehat{S}_{0}$ be the strict transform of $S_{0}$ and $D$ be the exceptional divisor of $\pi$. Let $E_{0}:=\widehat{S}_{0}\cap D$. Then $E_{0}=\mathbb{P}(\mathcal{N}^{*}_{C_{0}/S_{0}})$. Note that as ${\rm codim}_{C}(C\setminus C_{0})\geq 2$, we can extend $\mathcal{N}_{C_{0}/S_{0}}$ to be a reflexive subsheaf $\mathcal{F}\subset T_{C/X}$ on $C$. Then $E_{0}$ is a Zariski open subset of the compact set $\mathbb{P}(\mathcal{F}^{*})$.
Set $a:=\nu(\varphi|_{S_{0}},C_{0})$. We consider the current $T:=\pi^{*}(\omega_{X}+i\partial\overline{\partial}\varphi)\geq 0$ on $\widehat{X}$. By construction, the restriction $T|_{\widehat{S}_{0}}$ is well defined. By Proposition 2.1, we have
$a=\nu(T|_{\widehat{S}_{0}},E_{0}).$
Now we set $T_{a}:=T|_{\widehat{S}_{0}}-a[E_{0}]$. Then $T_{a}$ is still a positive current on $\widehat{S}_{0}$, and $T_{a}|_{E_{0}}$ is a well-defined positive current on $E_{0}$. Here we use the fact that $\varphi|_{S_{0}}$ has algebraic singularities.
Let $E$ be a desingularisation of $\mathbb{P}(\mathcal{F}^{*})$. Note that ${\rm codim}_{E}(E\setminus E_{0})$ might be 1, and the singularity of $T_{a}|_{E_{0}}$ might be transcendental near the boundary because of the transcendentality of $S_{0}$. Therefore, we cannot à priori extend $T_{a}|_{E_{0}}$ to be a current on $E$.
Let $\pi_{E_{0}}:E_{0}=\mathbb{P}(\mathcal{N}^{*}_{C_{0}/S_{0}})\to C_{0}$. Let $r$ be the dimension of the generic fiber of $\pi_{E_{0}}$, and let $\omega_{\widehat{X}}$ be a Kähler metric on $\widehat{X}$. After multiplying by some constant, we can assume that $(\pi_{E_{0}})_{*}\left((\omega_{\widehat{X}}|_{E_{0}})^{r}\right)=1$ at $y_{0}$. We consider the push-forward
$T_{C_{0}}:=(\pi_{E_{0}})_{*}\left(T_{a}|_{E_{0}}\wedge(\omega_{\widehat{X}}|_{E_{0}})^{r}\right).$
By construction, the $(1,1)$-current $T_{C_{0}}\geq 0$ on $C_{0}$. As ${\rm codim}_{C}(C\setminus C_{0})\geq 2$, by Proposition 2.1, we can extend it to be a positive current on $C$, denoted by $T_{C}$.
Thanks to Proposition 2.3, we have
$\nu(T_{C},y_{0})\geq\nu(T|_{\widehat{S}_{0}},\pi_{E_{0}}^{-1}(y_{0}))-a=\nu(\varphi|_{S_{0}},y_{0})-a,$
where the second equality comes from Proposition 2.1. Note that by Lemma 3.3 below, the cohomology class of $T_{C}$ is bounded by $(aK_{1}+K_{2})\cdot c_{1}(\omega_{X}|_{C})$ for some uniform constants $K_{1}$ and $K_{2}$. Then $\nu(T_{C},y_{0})$ is uniformly upper-bounded by $K\cdot(aK_{1}+K_{2})$, where $K$ is a constant depending on $c_{1}(\omega_{X}|_{C})$ and the Seshadri constant of $y_{0}$ in $C$. Therefore, we have
$K\cdot(aK_{1}+K_{2})\geq\nu(T_{C},y_{0})\geq\nu(\varphi|_{S_{0}},y_{0})-a,$
and the proposition is proved. ∎
Lemma 3.3.
Let $T_{C}$ be the current on $C$ constructed in the proof of Proposition 3.2. Then there exists a uniform constants $K_{1}$ and $K_{2}$ independent of $a$, such that $(aK_{1}+K_{2})\cdot c_{1}(\omega_{X}|_{C})-T_{C}$ is a psef class.
Proof.
Let $\alpha:=c_{1}(\pi^{*}\omega_{X})-a\cdot c_{1}(\mathcal{O}_{\widehat{X}}(D))\in H^{1,1}(\widehat{X},\mathbb{R})$. Let $\rho:E\to\mathbb{P}(\mathcal{F}^{*})$ be a desingularization and $\pi_{E}:E\to C$ be the projection. Consider the class
$\beta:=(\pi_{E})_{*}\Big{(}\rho^{*}\big{(}\big{(}\alpha\wedge(\omega_{\widehat{X}})^{r}\big{)}|_{\mathbb{P}(\mathcal{F}^{*})}\big{)}\Big{)}.$
Note that $\beta$ is a $(1,1)$-class on $C$ and is upper bounded by $(aK_{1}+K_{2})\cdot c_{1}(\omega_{X}|_{C})$ for some uniform constants $K_{1}$ and $K_{2}$ independent of $a$. To prove the lemma, it suffices to show that
(2) $c_{1}(\beta)=c_{1}(T_{C})\qquad\text{on }C.$
Let $s_{D}\in H^{0}(\widehat{X},\mathcal{O}_{\widehat{X}}(D))$ be the canonical section (which vanishes along $D$), and fix a smooth metric $h_{D}$ on $\mathcal{O}_{\widehat{X}}(D)$. Consider the current
$T_{\widehat{X}}:=\pi^{*}\omega_{X}+i\partial\overline{\partial}\varphi-ai\partial\overline{\partial}\log|s_{D}|^{2}_{h_{D}}-ai\Theta_{h_{D}}(\mathcal{O}_{\widehat{X}}(D))\qquad\text{on }\widehat{X}.$
Then we have $c_{1}(T_{\widehat{X}})=\alpha$ and $T_{\widehat{X}}|_{\widehat{S}_{0}}=T_{a}$ on $\widehat{S}_{0}$.
Set $A:=\pi^{*}\omega_{X}-ai\Theta_{h_{D}}(\mathcal{O}_{\widehat{X}}(D))$. Then $A$ is a smooth form in the class of $\alpha$, and we have
$A|_{E_{0}}=T_{a}|_{E_{0}}+i\partial\overline{\partial}\psi\qquad\text{on }E_{0},$
where $\psi=((\varphi-a\log|s_{D}|^{2}_{h_{D}})|_{\widehat{S}_{0}})|_{E_{0}}$ is well-defined on $E_{0}$. As a consequence,
$A|_{E_{0}}\wedge\big{(}\omega_{\widehat{X}}|_{E}\big{)}^{r}=T_{a}|_{E_{0}}\wedge\big{(}\omega_{\widehat{X}}|_{E}\big{)}^{r}+i\partial\overline{\partial}(\psi\cdot\big{(}\omega_{\widehat{X}}|_{E}\big{)}^{r})\qquad\text{on }E_{0}.$
Therefore, we have
(3) $p_{*}(A|_{E_{0}}\wedge\big{(}\omega_{\widehat{X}}|_{E}\big{)}^{r})=T_{C_{0}}+i\partial\overline{\partial}\Phi\qquad\text{on }C_{0}$
for some function $\Phi$ defined on $C_{0}$. Since ${\rm codim}_{C}(C\setminus C_{0})\geq 2$, by Proposition 2.1, the equality (3) extends to the whole space $C$. Note that the class of the extended current on the left-hand side of (3) is precisely $\beta$. This completes the proof of (2). ∎
Now we are going to prove Theorem 1.1. In the setting of Theorem 1.1, we fix a point $y_{0}\in C_{0}$. If $M$ is smooth, then by applying Proposition 3.1 to $(M,S_{0},y_{0})$, we can construct a $\omega_{X}$-psh function $\varphi$ on $M$ such that $\nu(\varphi|_{S_{0}},y_{0})$ is arbitrarily large. By applying the extension theorem [CT14, Thm 1.1], we obtain a $\omega_{X}$-psh function $\varphi$ on $X$ such that $\nu(\varphi|_{S_{0}},y_{0})$ is arbitrarily large. Together with Proposition 3.2, we conclude that $\nu(\varphi|_{S_{0}},C_{0})$ is arbitrarily large. This implies that $\mathcal{N}^{*}_{C_{0}/S_{0}}$ is psef (cf. the proof of Theorem 1.1). In the general case, when $M$ is not necessarily smooth, we have the following proposition.
Proposition 3.4.
In the setting of Theorem 1.1, let $\pi:\widehat{X}\to X$ be the blow-up of $C$ in $X$. Let $\widehat{S}_{0}$ be the strict transform of $S_{0}$, and let $D$ be the exceptional divisor of $\pi$. Define $E_{0}:=\widehat{S}_{0}\cap D$. Let $\omega_{\widehat{X}}$ be a Kähler metric on $\widehat{X}$.
If $\dim M>\dim S_{0}$, then for any $m\in\mathbb{N}$, there exists a $\omega_{\widehat{X}}$-psh function $\varphi_{m}$ on $\widehat{X}$ such that $\varphi_{m}|_{\widehat{S}_{0}}$ has analytic singularities and $\nu(\varphi_{m}|_{\widehat{S}_{0}},E_{0})\geq m$.
Proof.
Let $\widehat{M}\subset\widehat{X}$ be the strict transform of $M$. We desingularize the singular locus of $\widehat{M}$ and obtain a bimeromorphism $p:\widetilde{X}\to\widehat{X}$. Let $\widehat{S}_{0}$ be the strict transform of $S_{0}$ in $\widehat{X}$. Since $S_{0}$ is dense in $M$, $\widehat{S}_{0}$ is not contained in the singular locus of $\widehat{M}$. Let $\widetilde{S}_{0}\subset\widetilde{X}$ be the strict transform of $\widehat{S}_{0}$. Define $\widetilde{E}_{0}\subset\widetilde{S}_{0}$ as the unique component of $p^{-1}(E_{0})\cap\widetilde{S}_{0}$ such that $p(\widetilde{E}_{0})=E_{0}$. The uniqueness follows from the fact that $E_{0}$ has codimension $1$ in $\widehat{S}_{0}$ and that $\widehat{S}_{0}$ is smooth. We have the following diagram:
[…]
Let $\widetilde{y}_{0}\in\widetilde{E}_{0}$ be a generic point. Since $E_{0}\subset\widehat{S}_{0}$ has codimension one, the bimeromorphism $p:\widetilde{S}_{0}\to\widehat{S}_{0}$ is an isomorphism in a neighborhood of the generic point $\widetilde{y}_{0}$. In particular, $\widetilde{S}_{0}$ is smooth in a neighborhood of $\widetilde{y}_{0}$. Set $\widehat{y}_{0}:=p(\widetilde{y}_{0})$ and $y_{0}:=\pi(\widehat{y}_{0})$. Now, we fix a Kähler metric $\omega_{X}$ on $X$. By using [DP04], there exist Kähler metrics $\omega_{\widetilde{X}}$ on $\widetilde{X}$ and $\omega_{\widehat{X}}$ on $\widehat{X}$ such that
$p_{*}(c_{1}(\omega_{\widetilde{X}}))=c_{1}(\omega_{\widehat{X}})\quad\text{and}\quad\pi_{*}(c_{1}(\omega_{\widehat{X}}))=c_{1}(\omega_{X}).$
Since $\widetilde{M}$ is smooth and $\widetilde{S}_{0}$ is smooth in a neighborhood of $\widetilde{y}_{0}$, we can apply Proposition 3.1 to $(\widetilde{M},\widetilde{S}_{0},\widetilde{y}_{0})$. Together with the extension theorem [CT14, Thm 1.1], for any $m>0$, we can find a $\omega_{\widetilde{X}}$-psh function $\widetilde{\varphi}_{m}$ on $\widetilde{X}$ such that $\widetilde{\varphi}_{m}|_{\widetilde{S}_{0}}$ has analytic singularities and
$\nu(\widetilde{\varphi}_{m}|_{\widetilde{S}_{0}},\widetilde{y}_{0})\geq m.$
Set $\widetilde{T}_{m}:=\omega_{\widetilde{X}}+i\partial\overline{\partial}\widetilde{\varphi}_{m}\geq 0$. Then
$\nu(\widetilde{T}_{m}|_{\widetilde{S}_{0}},\widetilde{y}_{0})\geq m.$
We consider the push-forward
$\widehat{T}_{m}:=p_{*}(\widetilde{T}_{m}),\quad T_{m}:=(p\circ\pi)_{*}\widetilde{T}_{m}.$
Then
$c_{1}(\widehat{T}_{m})=c_{1}(\omega_{\widehat{X}}),\quad c_{1}(T_{m})=c_{1}(\omega_{X}),$
and the restrictions $\widehat{T}_{m}|_{\widehat{S}_{0}}$ and $T_{m}|_{S_{0}}$ are well defined. By using Lemma 3.5, there exists a constant $C_{1}$ independent of $m$ such that
$\left|\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},\widehat{y}_{0})-\nu(\widetilde{T}_{m}|_{\widetilde{S}_{0}},\widetilde{y}_{0})\right|\leq C_{1}.$
Therefore, we have
$\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},\widehat{y}_{0})\geq m-C_{1}.$
There are two possibilities:
- If $\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},E_{0})\geq\frac{m}{2}$ for infinitely many $m\in\mathbb{N}^{*}$, then the proposition is proved.
- If $\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},E_{0})\leq\frac{m}{2}$ for every sufficiently large $m$, then, thanks to Lemma 2.2, we have
$\nu(T_{m}|_{S_{0}},y_{0})\geq\frac{1}{K}\cdot\left(\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},\widehat{y}_{0})-\nu(\widehat{T}_{m}|_{\widehat{S}_{0}},E_{0})\right)\geq\frac{1}{K}\cdot\left(\frac{m}{2}-C_{1}\right)$
for some uniform constant $K$ independent of $m$. By applying Proposition 3.2, we obtain
$\nu(T_{m}|_{S_{0}},C_{0})\geq\frac{\nu(T_{m}|_{S_{0}},y_{0})-K_{2}}{K_{1}}\geq\widetilde{K}_{1}m-\widetilde{K}_{2}.$
for some uniform constants $\widetilde{K}_{1}>0$ and $\widetilde{K}_{2}$ independent of $m$. We write $T_{m}=\omega_{X}+i\partial\overline{\partial}\varphi_{m}$ on $X$. Then $\pi^{*}\varphi_{m}$ satisfies the proposition. ∎
Lemma 3.5.
There exists a constant $C_{1}$ independent of $m$ such that
$|\nu(\widehat{T}|_{\widehat{S}_{0}},\widehat{y}_{0})-\nu(\widetilde{T}|_{\widetilde{S}_{0}},\widetilde{y}_{0})|\leq C_{1}.$
Proof.
Let $\sum D_{i}$ be the exceptional divisor of $p$. By construction, we have
$c_{1}(\omega_{\widetilde{X}})=p^{*}c_{1}(\omega_{\widehat{X}})+\sum a_{i}[D_{i}]$
for some constants $a_{i}$. Then, we can write $\widetilde{T}=\omega_{\widetilde{X}}+i\partial\overline{\partial}\widetilde{\varphi}$ and $\widehat{T}=\omega_{\widehat{X}}+i\partial\overline{\partial}\psi$ for some function $\psi$ on $\widehat{X}$. Since $c_{1}(p_{*}(\omega_{\widetilde{X}}))=c_{1}(\omega_{\widehat{X}})$, there exists a function $\Phi$ on $\widehat{X}$ such that
$p_{*}(\omega_{\widetilde{X}})=\omega_{\widehat{X}}+i\partial\overline{\partial}\Phi.$
As a consequence, we have the pointwise identity
$p_{*}(\widetilde{\varphi})=\psi-\Phi+C_{0}\qquad\text{on }\widehat{X}$
for some constant $C_{0}$.
Since $p|_{\widetilde{S}_{0}}$ is an isomorphism near $\widetilde{y}_{0}$, we can find a neighborhood $\widetilde{U}\subset\widetilde{S}_{0}$ of $\widetilde{y}_{0}$ and a neighborhood $\widehat{U}\subset\widehat{S}_{0}$ of $\widehat{y}_{0}$ such that $p|_{\widetilde{U}}:\widetilde{U}\to\widehat{U}$ is an isomorphism. Consequently, at the level of quasi-psh functions, we have the identity
$(\psi-\Phi+C_{0})\circ(p|_{\widetilde{U}})=\widetilde{\varphi}\qquad\text{on }\widetilde{U}.$
Thus, we obtain
$\nu((\psi-\Phi+C_{0})|_{\widehat{S}_{0}},\widehat{y}_{0})=\nu(\widetilde{\varphi}|_{\widetilde{S}_{0}},\widetilde{y}_{0}).$
Since $\Phi$ depends only on the blow-up and the coefficients $a_{i}$, and is well-defined on $\widehat{S}_{0}$, we conclude that
$|\nu(\psi|_{\widehat{S}_{0}},\widehat{y}_{0})-\nu(\widetilde{\varphi}|_{\widetilde{S}_{0}},\widetilde{y}_{0})|\leq C_{1}$
for some uniform constant $C_{1}$. ∎
Now we can prove the main theorem 1.1:
Proof of the main theorem 1.1.
Let $\pi:\widehat{X}\to X$ be the blow-up of $C$ in $X$. Let $\widehat{S}_{0}$ be the strict transform of $S_{0}$, and let $D$ be the exceptional divisor of $\pi$. Define $E_{0}:=\widehat{S}_{0}\cap D$. Then we have $E_{0}=\mathbb{P}(\mathcal{N}^{*}_{C_{0}/S_{0}})$.
By applying Proposition 3.4, for every $m\in\mathbb{N}$, there exists a $\omega_{\widehat{X}}$-psh function $\varphi_{m}$ such that $a_{m}:=\nu(\varphi_{m}|_{\widehat{S}_{0}},E_{0})\geq m$, where $\omega_{\widehat{X}}$ is a Kähler metric on $\widehat{X}$.
Consider the current $T_{m}:=\omega_{\widehat{X}}+i\partial\overline{\partial}\varphi_{m}\geq 0$ on $\widehat{X}$. By construction, the restriction $T_{m}|_{\widehat{S}_{0}}$ is well defined. Now, we set
$G_{m}:=T_{m}|_{\widehat{S}_{0}}-a_{m}[E_{0}].$
Then, $G_{m}|_{E_{0}}$ is a well-defined positive current on $E_{0}$, and we have
(4) $c_{1}(G_{m}|_{E_{0}})=c_{1}(\omega_{\widehat{X}}|_{E_{0}}+a_{m}\cdot\mathcal{O}_{\mathbb{P}(\mathcal{N}^{*}_{C_{0}/S_{0}})}(1))\qquad\text{on }E_{0}.$
Since $a_{m}\geq m$ and $G_{m}|_{E_{0}}\geq 0$, we conclude that $\mathcal{N}^{*}_{C_{0}/S_{0}}$ is psef in the sense of Definition 2.4. This completes the proof of the theorem. ∎